Check whether the intervals [2,5] and [7,10] are equivalent or not


1
Expert's answer
2021-03-30T16:24:26-0400

Consider the function "f:[2,5]\to[7,10],\ f(x)=x+5." For each "y\in[7,10]" the equation "f(x)=y" (which is equivalent to "x+5=y") has a unique solution "x=y-5\in[2,5]". Therefore, the function "f" is a bijection. It follows that the intervals "[2,5]" and "[7,10]" are equivalent.